OpenStudy (anonymous):

Annabelle and Navene are multiplying (4275)(4372). Annabelle's Work Navene's Work (4275)(4372) = 42 + 375 + 2 = 4577 (4275)(4372) = 42⋅375⋅2 = 46710 Is either of them correct? Explain your reasoning.

OpenStudy (anonymous):

(4^2 7^5) (4^3 7^2)= 4^2 + ^3 7^5+ ^2 = 4^5 7^7

OpenStudy (anonymous):

annabele

OpenStudy (anonymous):

now ima do navenes

OpenStudy (freckles):

what do the carets before the numbers mean?

OpenStudy (freckles):

for example what does ^3 mean?

OpenStudy (anonymous):

(4^2 7^5) (4^3 7^2) = 4^2 x ^3 7^5 x ^2 = 4^6 7^10

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

@freckles it means like its an exponent so i guess its raised

OpenStudy (michele_laino):

hint:we can write this: \[\large \left( {{4^2} \cdot {7^5}} \right)\left( {{4^3} \cdot {7^2}} \right) = \left( {{4^2} \cdot {4^3}} \right)\left( {{7^5} \cdot {7^2}} \right) = {4^{2 + 3}} \cdot {7^{5 + 2}} = ...?\]

OpenStudy (freckles):

like I know what 5^2 means this is 5*5 but I have no clue what something like ^3 means

OpenStudy (anonymous):

the exponents are being added or also they are being multiplied in the other one

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @77777jeannie77777 (4^2 7^5) (4^3 7^2)= 4^2 + ^3 7^5+ ^2 = 4^5 7^7 \(\color{blue}{\text{End of Quote}}\) you ate base

OpenStudy (anonymous):

whats that?

OpenStudy (anonymous):

can u see it?

OpenStudy (anonymous):

its #7

Nnesha (nnesha):

http://prntscr.com/89limo here what number is raising to the 3rd power that's what she is asking

OpenStudy (anonymous):

can u see the question?

Nnesha (nnesha):

no :(

OpenStudy (anonymous):

ugh

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @77777jeannie77777 (4^2 7^5) (4^3 7^2)= 4^2 + ^3 7^5+ ^2 = 4^5 7^7 \(\color{blue}{\text{End of Quote}}\) it's 4^3 power not 4^2 ( i guess )

OpenStudy (anonymous):

wait a sec im going to see if i can make it a file

OpenStudy (anonymous):

no i cant :(

OpenStudy (anonymous):

how do i take a screenshot?

Nnesha (nnesha):

okay just recheck ur work #first comment

OpenStudy (anonymous):

how do i take one?

OpenStudy (anonymous):

ok wait a sec plz

OpenStudy (anonymous):

@Michele_Laino @freckles

Nnesha (nnesha):

okay so 4^{3+2} use parentheses so we can understand otay :=)

OpenStudy (anonymous):

ok:D

OpenStudy (anonymous):

4^(5)

OpenStudy (michele_laino):

as I wrote before, your expression can be rewritten as below: \[\large \left( {{4^2} \cdot {7^5}} \right)\left( {{4^3} \cdot {7^2}} \right) = \left( {{4^2} \cdot {4^3}} \right)\left( {{7^5} \cdot {7^2}} \right) = {4^{2 + 3}} \cdot {7^{5 + 2}} = ...?\]

Nnesha (nnesha):

when we multiply same bases we should add or multiply the exponents ?

OpenStudy (anonymous):

add?

Nnesha (nnesha):

yes right

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ok so michele its 4^(5) times 7^(7)

OpenStudy (anonymous):

right?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

so then when there not like bases how do u multiply it?

OpenStudy (michele_laino):

we have to compute the explicit value of each power and then we have to multiply them together

OpenStudy (anonymous):

whaa?

Nnesha (nnesha):

then you can't like for this one \[2^2 \times 3^5 \]

Nnesha (nnesha):

you will not see question like that one ^^ hehe

OpenStudy (anonymous):

lol

OpenStudy (michele_laino):

for example: \[\large \begin{gathered} \left( {{4^2} \cdot {7^5}} \right)\left( {{4^3} \cdot {7^2}} \right) = \left( {{4^2} \cdot {4^3}} \right)\left( {{7^5} \cdot {7^2}} \right) = {4^{2 + 3}} \cdot {7^{5 + 2}} = \hfill \\ \hfill \\ = 1024 \cdot 823543 = {\text{843}}{\text{,308}}{\text{,032}} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ok nnesha i didnt understand what u said above that

OpenStudy (anonymous):

whoa!!

OpenStudy (anonymous):

so u just multiply the bases and then the exponents together right?

Nnesha (nnesha):

you question is when bases are not like then how would we multiply it ? if bases are not same you can't multiply or add the exponents \[2^2 \times 3^5= 2 \times 2 \times 3 \times 3 \times 3 \times 3 \times 3\]

OpenStudy (michele_laino):

no, I'm sorry, you have to evaluate each power first, then you have to multiply the obtained values together

OpenStudy (anonymous):

ohh then you multiply all those together?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ooh

Nnesha (nnesha):

btw ur username is same as \[\huge\rm 7^5jeannie7^5\]

OpenStudy (anonymous):

haha ^.^

OpenStudy (anonymous):

so 972?

Nnesha (nnesha):

o^_^o

Nnesha (nnesha):

what 972 ?

OpenStudy (anonymous):

multiplying all of them together

Nnesha (nnesha):

ohh the example i gave you ? yes

OpenStudy (anonymous):

can you please do a summarry of the answer please? of the question lol

Nnesha (nnesha):

you already know when we multiply same bases we should `add or multiply` their exponents ?

OpenStudy (anonymous):

add

Nnesha (nnesha):

yes! right! so who is right anna or Navene ?

OpenStudy (anonymous):

ohh annabel

Nnesha (nnesha):

yes yayay! anna!

OpenStudy (anonymous):

yayyyyy thx

Nnesha (nnesha):

my pleasure :=) , \[\huge\rm 7^5jeannie7^5\]

OpenStudy (anonymous):

;D